# Orlicz-Sobolev Transport for Unbalanced Measures on a Graph

### v1

@@ Abstract
Moving beyond $L^{p}$ geometric structure, Orlicz-Wasserstein (OW): equal contribution leverages a specific class of convex functions for Orlicz geometric structure. While OW remarkably helps to advance certain machine learning approaches, it has a high computational complexity due to its two-level optimization formula. Recently, Le et al. (2024) exploits graph structure to propose generalized Sobolev transport (GST), i.e., a scalable variant for OW. However, GST assumes that input measures have the same mass. Unlike optimal transport (OT), it is nontrivial to incorporate a mass constraint to extend GST for measures on a graph, possibly having different total mass. In this work, we propose to take a step back by considering the entropy partial transport (EPT) for nonnegative measures on a graph. By leveraging Caffarelli & McCann (2010)’s observations, EPT can be reformulated as a standard complete OT between two corresponding balanced measures. Consequently, we develop a novel EPT with Orlicz geometric structure, namely Orlicz-EPT, for unbalanced measures on a graph. Especially, by exploiting the dual EPT formulation and geometric structures of the graph-based Orlicz-Sobolev space, we derive a novel regularization to propose Orlicz-Sobolev transport (OST). The resulting distance can be efficiently computed by simply solving a univariate optimization problem, unlike the high-computational two-level optimization problem for Orlicz-EPT. Additionally, we derive geometric structures for the OST and draw its relations to other transport distances. We empirically show that OST is several-order faster than Orlicz-EPT. We further illustrate preliminary evidences on the advantages of OST for document classification, and several tasks in topological data analysis.
@@ 1 Introduction
Moving beyond $L^{p}$ geometric structure, Orlicz-Wasserstein (OW) leverages a specific class of convex functions for Orlicz geometric structure. Intuitively, OW is an instance of optimal transport (OT), which utilizes Orlicz metric as its ground cost (Sturm, 2011; Kell, 2017; Guha et al., 2023; Altschuler & Chewi, 2023; Le et al., 2024). Notably, OW remarkably helps to advance certain machine learning approaches. For examples, Altschuler & Chewi (2023) recently leverage OW as a metric shift for Rényi divergence and propose novel differential-privacy-inspired techniques to overcome longstanding challenges for proving fast convergence of hypocoercive differential equations. Additionally, Guha et al. (2023) leverage OW metric to alleviate several raised concerns caused from the usage of classical OT with Euclidean ground cost for quantifying the rates of parameter convergence within infinite Gaussian mixtures to significantly improve the Bayesian contraction rate of parameters arising from hierarchical Bayesian nonparametric models. However, OW has a high computational complexity steaming from its two-level optimization problem, i.e., one level for optimization plan, and the other level for an extra scalar in Orlicz metric structure. Recently, Le et al. (2024) propose generalized Sobolev transport, which is a scalable variant of OW for probability measures on a graph, for practical applications, especially for large-scale settings. Besides that, Orlicz geometric structure has been also applied for several machine learning problems, e.g., linear regression (Andoni et al., 2018; Song et al., 2019), scalable approaches for reinforcement learning, kernelized support vector machine, and clustering (Deng et al., 2022). Orlicz metric is also used to derive a finite-sample deviation bound for a general class of polynomial-growth functions to approximate high-order derivatives for arbitrary kernel (Chamakh et al., 2020), and as an OT regularization (Lorenz & Mahler, 2022).
When input measures have different total mass, several proposals have been developed in the literature (Hanin, 1992; Guittet, 2002; Benamou, 2003; Caffarelli & McCann, 2010; Figalli, 2010; Lellmann et al., 2014; Piccoli & Rossi, 2014, 2016; Frogner et al., 2015; Kondratyev et al., 2016; Liero et al., 2018; Chizat et al., 2018; Bonneel & Coeurjolly, 2019; Gangbo et al., 2019; Séjourné et al., 2019, 2022; Pham et al., 2020; Sato et al., 2020; Chapel et al., 2020; Balaji et al., 2020; Mukherjee et al., 2021; Le & Nguyen, 2021; Le et al., 2023; Nguyen et al., 2023), to name a few. The unbalanced setting for nonnegative measures has several applications, e.g., for color transfer and shape matching (Bonneel et al., 2015); multi-label learning (Frogner et al., 2015); positive-unlabeled learning (Chapel et al., 2020); natural language processing and topological data analysis (Le & Nguyen, 2021; Le et al., 2023); robust approaches for applications with noisy supports or outliers (Frogner et al., 2015; Balaji et al., 2020; Mukherjee et al., 2021).
In this work, we focus on OT problem with Orlicz geometric structure for unbalanced measures supported on a graph metric space (Le et al., 2022). Although GST (Le et al., 2023) provides a scalable variant for OW, GST assumes that input measures have the same mass as in OW. Unlike OT, it is nontrivial to incorporate a mass constraint to extend GST for general nonnegative measures on a graph. To address this challenge, our key insight is to take a step back to leverage Caffarelli & McCann (2010)’s observations, which allows to reformulate unbalanced optimal transport (UOT) problem as corresponding standard complete OT problem (Le et al., 2023), to propose Orlicz-EPT for unbalanced measures on a graph with Orlicz geometric structure. Furthermore, by exploiting the graph structure, we propose Orlicz-Sobolev transport, which scales Orlicz-EPT for practical applications.
Contribution. In summary, our contributions are two-fold as follows:
i) We propose to take a detour, and leverage EPT problem on a graph for unbalanced measures on a graph. By exploiting Caffarelli & McCann (2010)’s observations, we reformulate it as a corresponding standard complete OT to derive the proposed Orlicz-EPT. By further exploiting the graph structure, we develop a novel regularization and propose Orlicz-Sobolev transport (OST) which scales Orlicz-EPT for practical applications by showing that OST can be computed by solving a univariate optimization problem.
@@
ii) We derive theoretical results for OST and draw its connections to other transport distances. We empirically illustrate that OST is several-order faster than Orlicz-EPT, and show preliminary evidences on the advantages of OST for document classification and topological data analysis.
@@
We briefly review related notions for our proposals in §2. In §3, we describe our proposed Orlicz-EPT and Orlicz-Sobolev transport (OST) based on EPT problem for unbalanced measures on a graph. Then, we derive theoretical properties for OST and draw its connections to other transport distances in §4. In §5, we discuss relations of the proposed approaches with other transport distances in the literature. We provide experimental results in §6, and conclude our work in §7.
@@ 2 Preliminaries
Graph. We use the same graph setting as in (Le et al., 2022). Specifically, let $V$ and $E$ be respectively the sets of nodes and edges. We consider a connected, undirected, and physical [^1] In the sense that $V$ is a subset of Euclidean space $\mathbb{R}^{n}$, and each edge $e\in E$ is the standard line segment in $\mathbb{R}^{n}$ connecting the two vertices of the edge $e$. graph ${\mathbb{G}}=(V,E)$ with positive edge lengths $\{w_{e}\}_{e\in E}$. Following the convention in (Le et al., 2022) for continuous graph setting, ${\mathbb{G}}$ is regarded as the set of all nodes in $V$ together with all points forming the edges in $E$. Also, ${\mathbb{G}}$ is equipped with the graph metric $d_{{\mathbb{G}}}(x,y)$ which equals to the length of the shortest path in ${\mathbb{G}}$ between $x$ and $y$. Additionally, we assume that there exists a fixed root node $z_{0}\in V$ such that the shortest path connecting $z_{0}$ and $x$ is unique for any $x\in{\mathbb{G}}$, i.e., the uniqueness property of the shortest paths (Le et al., 2022).
Let $[x,z]$ denote the shortest path connecting $x$ and $z$ in ${\mathbb{G}}$. For $x\in{\mathbb{G}}$, edge $e\in E$, define the sets $\Lambda(x)$, $\gamma_{e}$ as follows:
$\Lambda(x):=\big{\{}y\in{\mathbb{G}}:\,x\in[z_{0},y]\big{\}},$
$\gamma_{e}:=\big{\{}y\in{\mathbb{G}}:\,e\subset[z_{0},y]\big{\}}.$ (1)
@@
Functions on graph. By a continuous function $f$ on ${\mathbb{G}}$, we mean that $f:{\mathbb{G}}\to\mathbb{R}$ is continuous w.r.t. the topology on ${\mathbb{G}}$ induced by the Euclidean distance. Henceforth, $C({\mathbb{G}})$ denotes the set of all continuous functions on ${\mathbb{G}}$. Similar notation is used for continuous functions on ${\mathbb{G}}\times{\mathbb{G}}$.
Given a scalar $b>0$, then a function $f:{\mathbb{G}}\to\mathbb{R}$ is called $b$-Lipschitz w.r.t. the graph metric $d_{\mathbb{G}}$ if
$|f(x)-f(y)|\leq b\,d_{\mathbb{G}}(x,y),\,\,\forall x,y\in{\mathbb{G}}.$
A family of convex functions. We consider the collection of $N$-functions (Adams & Fournier, 2003, §8.2) which are special convex functions on $\mathbb{R}_{+}$. Hereafter, a strictly increasing and convex function $\Phi:[0,\infty)\to[0,\infty)$ is called an $N$-function if $\lim_{t\to 0}\frac{\Phi(t)}{t}=0$ and $\lim_{t\to+\infty}\frac{\Phi(t)}{t}=+\infty$.
@@
3 Orlicz-Sobolev Transport
In this section, for completeness, we briefly review the entropy partial transport (EPT) problem for unbalanced measures on a graph (Le et al., 2023, §3). We extend it for Orlicz geometric structure. Then, we describe our proposed Orlicz-Sobolev transport.
3.1 Orlicz-EPT: EPT on a Graph with Orlicz Structure
EPT on a graph.
Let $\gamma_{1},\gamma_{2}$ respectively be the first and second marginals of $\gamma\in{\mathcal{P}}({\mathbb{G}}\times{\mathbb{G}})$. For unbalanced measures $\mu,\nu\in{\mathcal{P}}({\mathbb{G}})$, we consider the set $\Pi_{\leq}(\mu,\nu):=\left\{\gamma:\,\gamma_{1}\leq\mu,\,\gamma_{2}\leq\nu\right\}$ where $\gamma_{1}\leq\mu$ means $\gamma_{1}(B)\leq\mu(B)$ for all Borel set $B\subset{\mathbb{G}}$, similarly for $\gamma_{2}\leq\nu$. Additionally, let $f_{1},f_{2}$ respectively be the Radon-Nikodym derivatives of $\gamma_{1}$ w.r.t. $\mu$ and of $\gamma_{2}$ w.r.t. $\nu$, i.e., $\gamma_{1}=f_{1}\mu$ ($0\leq f_{1}\leq 1$, $\mu$-a.e.) and $\gamma_{2}=f_{2}\nu$ ($0\leq f_{2}\leq 1$, $\nu$-a.e.).
For convex and lower semicontinuous entropy functions $F_{1},\,F_{2}:[0,1]\to(0,\infty)$, and nonnegative weight functions $w_{1},w_{2}:{\mathbb{G}}\to[0,\infty)$, we consider the weighted relative entropies of $\gamma_{1}$ w.r.t. $\mu$ and of $\gamma_{2}$ w.r.t. $\nu$ as follows:
${\mathcal{F}}_{1}(\gamma_{1}|\mu)$ $:=\int_{\mathbb{G}}w_{1}(x)F_{1}(f_{1}(x))\mu(\mathrm{d}x),$ (3)
${\mathcal{F}}_{2}(\gamma_{2}|\nu)$ $:=\int_{\mathbb{G}}w_{2}(x)F_{2}(f_{2}(x))\nu(\mathrm{d}x).$ (4)
We consider the graph metric $d_{{\mathbb{G}}}$ for the ground cost function. For a positive scalar $b>0$, and scalar $m\in[0,\bar{m}]$ where $\bar{m}:=\min\{\mu({\mathbb{G}}),\nu({\mathbb{G}})\}$, the EPT problem on ${\mathbb{G}}$ is
$\mathrm{W}_{m}(\mu,\nu):=\inf_{\gamma\in\Pi_{\leq}(\mu,\nu),\,\gamma({\mathbb{G}}\times{\mathbb{G}})=m}\Big{[}{\mathcal{F}}_{1}(\gamma_{1}|\mu)+{\mathcal{F}}_{2}(\gamma_{2}|\nu)+\,b\int_{{\mathbb{G}}\times{\mathbb{G}}}d_{{\mathbb{G}}}(x,y)\gamma(\mathrm{d}x,\mathrm{d}y)\Big{]}.\hskip 1.99997pt$ (5)
As in Le et al. (2023, §3), by using the entropy functions $F_{1}(s)=F_{2}(s)=|s-1|$ and considering a Lagrange multiplier $\lambda\in\mathbb{R}$ conjugate to the constraint $\gamma({\mathbb{G}}\times{\mathbb{G}})=m$, we instead study the problem
@@ 3.1 Orlicz-EPT: EPT on a Graph with Orlicz Structure
where $\mathcal{C}_{\lambda}(\gamma)=\int_{\mathbb{G}}w_{1}\mu(\mathrm{d}x)+\int_{\mathbb{G}}w_{2}\nu(\mathrm{d}x)-\int_{\mathbb{G}}w_{1}\gamma_{1}(\mathrm{d}x)-\int_{\mathbb{G}}w_{2}\gamma_{2}(\mathrm{d}x)+b\int_{{\mathbb{G}}\times{\mathbb{G}}}[d_{{\mathbb{G}}}(x,y)-\lambda]\gamma(\mathrm{d}x,\mathrm{d}y)$.
By leveraging the observation of Caffarelli & McCann (2010), we reformulate problem (6) as the standard complete OT problem. More specifically, we consider a point $\hat{s}$ outside graph ${\mathbb{G}}$ (i.e., $\hat{s}\notin{\mathbb{G}}$). We then extend the graph metric ground cost $d_{{\mathbb{G}}}$ to a new cost function $\hat{c}$ with $b\lambda$-deviation for its nonnegativity on the set $\hat{\mathbb{G}}:={\mathbb{G}}\cup\{\hat{s}\}$ as follows:
@@
For unbalanced measures $\mu,\nu$, by adding a Dirac mass at point $\hat{s}$, we construct corresponding balanced unit-mass measures $\hat{\mu}=\frac{\mu+\nu({\mathbb{G}})\delta_{\hat{s}}}{\mu({\mathbb{G}})+\nu({\mathbb{G}})}$ and $\hat{\nu}=\frac{\nu+\mu({\mathbb{G}})\delta_{\hat{s}}}{\mu({\mathbb{G}})+\nu({\mathbb{G}})}$. Consequently, we consider the standard complete OT problem between $\hat{\mu},\hat{\nu}$ with cost $\hat{c}$ as follows
$\mathrm{KT}(\mu,\nu):=\left(\mu({\mathbb{G}})+\nu({\mathbb{G}})\right)\left(\mathcal{W}_{\hat{c}}(\hat{\mu},\hat{\nu})-b\lambda\right),$ (8)
where $\mathcal{W}_{\hat{c}}(\hat{\mu},\hat{\nu}):=\inf_{\tilde{\gamma}\in\Pi(\hat{\mu},\hat{\nu})}\int_{\hat{\mathbb{G}}\times\hat{\mathbb{G}}}\hat{c}(x,y)\tilde{\gamma}(\mathrm{d}x,\mathrm{d}y)$, and $\Pi(\hat{\mu},\hat{\nu}):=\Big{\{}\hat{\gamma}\in{\mathcal{P}}(\hat{\mathbb{G}}\times\hat{\mathbb{G}}):\hat{\mu}(U)=\hat{\gamma}(U\times\hat{\mathbb{G}}),\,\hat{\nu}(U)=\hat{\gamma}(\hat{\mathbb{G}}\times U)\mbox{ for all Borel sets }U\subset\hat{\mathbb{G}}\Big{\}}$.
Therefore, one can reformulate an unbalanced OT problem (6) into a corresponding standard complete OT (8). Consequently, we not only bypass all the technical challenges in the unbalanced setting but also utilize abundant existing results and approaches in the standard balanced setting for OT problem with unbalanced measures on a graph.
@@
The nonnegativity of the cost function $\hat{c}$ (7) for standard OT (6) is essential for the development of EPT with Orlicz geometric structure. We remark that such nonnegativity is not guaranteed in Le et al. (2023).
Orlicz-EPT.
Following the approaches in Sturm (2011); Kell (2017); Guha et al. (2023); Chewi (2023), we define Orlicz-EPT, i.e., EPT with Orlicz geometric structure, upon the corresponding standard OT problem (8) as follows:

### v2

# An Efficient Orlicz-Sobolev Approach for Transporting Unbalanced Measures on a Graph

@@ Abstract
We investigate optimal transport (OT) for measures on graph metric spaces with different total masses. To mitigate the limitations of traditional $L^{p}$ geometry, Orlicz-Wasserstein (OW) and generalized Sobolev transport (GST) employ Orlicz geometric structure, leveraging convex functions to capture nuanced geometric relationships and remarkably contribute to advance certain machine learning approaches. However, both OW and GST are restricted to measures with equal total mass, limiting their applicability to real-world scenarios where mass variation is common, and input measures may have noisy supports, or outliers. To address unbalanced measures, OW can either incorporate mass constraints or marginal discrepancy penalization, but this leads to a more complex two-level optimization problem. Additionally, GST provides a scalable yet rigid framework, which poses significant challenges to extend GST to accommodate nonnegative measures. To tackle these challenges, in this work we revisit the entropy partial transport (EPT) problem. By exploiting Caffarelli & McCann [12]’s insights, we develop a novel variant of EPT endowed with Orlicz geometric structure, called Orlicz-EPT. We establish theoretical background to solve Orlicz-EPT using a binary search algorithmic approach. Especially, by leveraging the dual EPT and the underlying graph structure, we formulate a novel regularization approach that leads to the proposed Orlicz-Sobolev transport (OST). Notably, we demonstrate that OST can be efficiently computed by simply solving a univariate optimization problem, in stark contrast to the intensive computation needed for Orlicz-EPT. Building on this, we derive geometric structures for OST and draw its connections to other transport distances. We empirically illustrate that OST is several-order faster than Orlicz-EPT. Furthermore, we show initial evidence on the advantages of OST for measures on a graph in document classification and topological data analysis.
@@ 1 Introduction
Orlicz-Wasserstein (OW) extends $L^{p}$ geometry by leveraging a specific class of convex functions for Orlicz geometric structure. Intuitively, OW is an instance of optimal transport (OT), which utilizes Orlicz metric as its ground cost [67, 32, 28, 3, 41]. Building on this foundation, OW has proven instrumental in advancing certain machine learning approaches. For example, recent works have leveraged OW to tackle challenging problems: Altschuler & Chewi [3] use OW as a metric shift for Rényi divergence, enabling novel differential-privacy-inspired techniques to overcome longstanding challenges for fast convergence of hypocoercive differential equations, while Guha et al. [28] employ OW metric to significantly improve Bayesian contraction rates in hierarchical Bayesian nonparametric models by overcoming limitations raised from the usage of traditional OT with Euclidean ground cost. However, OW’s high computational complexity, stemming from its two-level optimization formula, poses a significant limitation. To address this challenge, Le et al. [41] introduce generalized Sobolev transport (GST), a scalable variant of OW suitable for practical application domains, especially for large-scale settings. Moreover, Orlicz geometric structure has been successfully applied to various machine learning problems, including linear regression [4, 66], scalable approaches [21] for reinforcement learning, kernelized support vector machines, and clustering. Additionally, Orlicz metrics play a crucial role in deriving deviation bounds for polynomial-growth functions to approximate kernel derivatives [13], and have been used as regularization in OT problems [46]. For in-depth studies on Orlicz functions, see [2, 60].
When dealing with input measures having different total masses, various approaches have been proposed in the literature to address this challenge [30, 29, 6, 12, 25, 44, 57, 58, 26, 33, 45, 18, 8, 27, 63, 64, 56, 62, 15, 5, 48, 36, 24, 16, 65, 40, 52, 7, 14, 72]. These approaches for unbalanced measures have proven effective in various domains, including color transfer [8], shape matching [8], image-to-image translation [72], multi-label learning [26], positive-unlabeled learning [15], point-cloud gradient flow [72], natural language processing [36, 40], topological data analysis (TDA) [36, 40], generative modeling [5, 72], domain adaptation [5], and robust approaches for handling noisy supports, outliers [26, 5, 48], or noisy ground cost [54, 42].
In this work, we focus on the OT problem with Orlicz geometric structure for unbalanced measures supported on a graph metric space. On one hand, OW naturally extends OT’s flexibility to handle unbalanced measures by incorporating either mass constraints or marginal difference penalization, formulated as partial OT (POT) or unbalanced OT (UOT) respectively. However, these approaches result in a more complex two-level optimization problem, analogous to POT/UOT with Orlicz metric cost, which poses significant computational challenges. On the other hand, although GST provides a scalable alternative to the computationally intensive OW, it still assumes equal-mass input measures. Moreover, due to GST’s definition as an optimization over the critic function, extending it to unbalanced measures is nontrivial. To address these limitations, we revisit the entropy partial transport (EPT) problem [36, 40, 72] and leverage insights from Caffarelli & McCann [12] to reformulate EPT as a standard complete OT problem. Then by carefully calibrating the corresponding ground cost for its nonnegativity, we propose Orlicz-EPT and establish a theoretical foundation for solving it by a binary search algorithmic approach. Furthermore, by exploiting the dual EPT and underlying graph structure, we introduce a novel regularization approach, leading to Orlicz-Sobolev transport (OST), which scales Orlicz-EPT for practical applications.
Contribution.
In summary, our contributions are two-fold as follows:
  • We revisit the EPT problem and leverage Caffarelli & McCann [12]’s insights to reformulate EPT as a standard complete OT, leading to the development of the proposed Orlicz-EPT. We establish its theoretical foundations, enabling a binary search algorithmic approach for its computation. Additionally, we develop a novel regularization approach, resulting in the proposed OST. We show that OST can be efficiently computed by simply solving a univariate optimization problem, unlike the computationally intensive Orlicz-EPT.
@@
  • We derive geometric structures for OST and establish its connections to other transport distances. Our empirical results demonstrate that OST is several-order faster than Orlicz-EPT. We also provide initial evidence on the advantages of OST for document classification and TDA.
@@
In §2, we briefly review relevant background and notions. We revisit EPT problem and propose Orlicz-EPT in §3. In §4, we introduce the computationally efficient OST. Then we derive geometric structures for OST and draw its connections to other transport distances in §5. In §6, we discuss related work. Empirical results are presented in §7, followed by concluding remarks in §8. Proofs of key theoretical results and additional materials are deferred to the Appendices. Furthermore, we have released code for our proposed approaches. [^1] The code repository is on https://github.com/lttam/OST_OrliczEPT.
@@ 2 Preliminaries
In this section, we introduce notations, and briefly review graph, and Orlicz functions.
@@
$\Lambda(x):=\big\{y\in{\mathbb{G}}:\,x\in[z_{0},y]\big\},\qquad\gamma_{e}:=\big\{y\in{\mathbb{G}}:\,e\subset[z_{0},y]\big\}.$ (1)
Functions on graph. By a continuous function $f$ on $\mathbb{G}$}, we mean that $f:{\mathbb{G}}\to\mathbb{R}$ is continuous w.r.t. the topology on $\mathbb{G}$} induced by the Euclidean distance. Henceforth, $C({\mathbb{G}})$ denotes the set of all continuous functions on $\mathbb{G}$}. Similar notation is used for continuous functions on ${\mathbb{G}}\times{\mathbb{G}}$. Given a positive scalar $b>0$, then a function $f:{\mathbb{G}}\to\mathbb{R}$ is called $b$-Lipschitz w.r.t. $d_{\mathbb{G}}$ if $|f(x)-f(y)|\leq b\,d_{\mathbb{G}}(x,y)$ for every $x$ and $y$ in $\mathbb{G}$}.
A family of convex functions. We consider the collection of $N$-functions [2, §8.2] which are special convex functions on $\mathbb{R}_{+}$. Hereafter, a strictly increasing and convex function $\Phi:[0,\infty)\to[0,\infty)$ is called an $N$-function if $\lim_{t\to 0}\frac{\Phi(t)}{t}=0$ and $\lim_{t\to+\infty}\frac{\Phi(t)}{t}=+\infty$.
@@
3 Orlicz-EPT: Entropy Partial Transport with Orlicz Geometric Structure
In this section, we revisit the entropy partial transport (EPT) problem [36, 40, 72], then develop Orlicz-EPT as a variant of EPT endowed with Orlicz geometric structure.
3.1 Entropy Partial Transport (EPT)
[…]
For convex and lower semicontinuous entropy functions $F_{1},\,F_{2}:[0,1]\to(0,\infty)$, and nonnegative weight functions $w_{1},w_{2}:{\mathbb{G}}\to[0,\infty)$, we consider the weighted relative entropies ${\mathcal{F}}_{1}(\gamma_{1}|\mu):=\int_{\mathbb{G}}w_{1}(x)F_{1}(f_{1}(x))\mu(\mathrm{d}x)$, and ${\mathcal{F}}_{2}(\gamma_{2}|\nu):=\int_{\mathbb{G}}w_{2}(x)F_{2}(f_{2}(x))\nu(\mathrm{d}x)$. For scalar $b>0$, scalar $m\in[0,\bar{m}]$ with $\bar{m}:=\min\{\mu({\mathbb{G}}),\nu({\mathbb{G}})\}$, and graph metric $d_{{\mathbb{G}}}$ as ground cost, the EPT problem is
$\mathrm{W}_{m}(\mu,\nu):=\inf_{\gamma\in\Pi_{\leq}(\mu,\nu),\,\gamma({\mathbb{G}}\times{\mathbb{G}})=m}\Big[{\mathcal{F}}_{1}(\gamma_{1}|\mu)+{\mathcal{F}}_{2}(\gamma_{2}|\nu)+\,b\int_{{\mathbb{G}}\times{\mathbb{G}}}d_{{\mathbb{G}}}(x,y)\gamma(\mathrm{d}x,\mathrm{d}y)\Big].$ (3)
Following [40, §3], by using entropy functions $F_{1}(s)=F_{2}(s):=|s-1|$ and considering a Lagrange multiplier $\lambda\in\mathbb{R}$ conjugate to the constraint $\gamma({\mathbb{G}}\times{\mathbb{G}})=m$, we instead study the problem
@@ 3.1 Entropy Partial Transport (EPT)
where $\mathcal{C}_{\lambda}(\gamma)=\int_{\mathbb{G}}w_{1}\mu(\mathrm{d}x)+\int_{\mathbb{G}}w_{2}\nu(\mathrm{d}x)-\int_{\mathbb{G}}w_{1}\gamma_{1}(\mathrm{d}x)-\int_{\mathbb{G}}w_{2}\gamma_{2}(\mathrm{d}x)+b\int_{{\mathbb{G}}\times{\mathbb{G}}}[d_{{\mathbb{G}}}(x,y)-\lambda]\gamma(\mathrm{d}x,\mathrm{d}y)$. [^4] The relationship between Problem (3) and Problem (4) is established in [40, Theorem A.1].
EPT as a standard OT. Following Caffarelli & McCann [12]’s insights, we can reformulate problem (4) as the standard complete OT problem. However, it is not guarantee that the corresponding standard OT has a nonnegative ground cost, e.g., see [12, 36, 40, 72]. Therefore, such OT reformulation may not be applicable to derive corresponding OW as in [67, 32, 28, 3, 41] since $N$-function is only defined for nonnegative domain (§2). Therefore, it is essential to carefully calibrate the ground cost of the corresponding standard OT of EPT to ensure its nonnegativity.
Precisely, following [40, Theorem 3.1], we henceforth consider $\lambda\geq 0$. [^5] The dual EPT result is the foundation for developing Orlicz-Sobolev transport in §4, where $\lambda$ is nonnegative. Then let $\hat{s}$ be a point outside graph $\mathbb{G}$}, i.e., $\hat{s}\notin{\mathbb{G}}$, and extend graph metric cost $d_{{\mathbb{G}}}$ on $\mathbb{G}$} to a new nonnegative cost function $\hat{c}$ with $b\lambda$-deviation on $\hat{\mathbb{G}}:={\mathbb{G}}\cup\{\hat{s}\}$ as follows:
@@
For unbalanced measures $\mu,\nu$, we construct corresponding probability (balanced) measures $\hat{\mu}=\frac{\mu+\nu({\mathbb{G}})\delta_{\hat{s}}}{\mu({\mathbb{G}})+\nu({\mathbb{G}})}$ and $\hat{\nu}=\frac{\nu+\mu({\mathbb{G}})\delta_{\hat{s}}}{\mu({\mathbb{G}})+\nu({\mathbb{G}})}$. Let $\Pi(\hat{\mu},\hat{\nu}):=\Big\{\hat{\gamma}\in{\mathcal{P}}(\hat{\mathbb{G}}\times\hat{\mathbb{G}}):\hat{\mu}(U)=\hat{\gamma}(U\times\hat{\mathbb{G}}),\,\hat{\nu}(U)=\hat{\gamma}(\hat{\mathbb{G}}\times U)\mbox{ for all Borel sets }U\subset\hat{\mathbb{G}}\Big\}$, then one can recast EPT (4) as a standard OT with cost $\hat{c}$.
Proposition 3.1.
Consider the standard OT $\mathcal{W}_{\hat{c}}$ between probability measures $\hat{\mu},\hat{\nu}$ with cost $\hat{c}$,
$\mathcal{W}_{\hat{c}}(\hat{\mu},\hat{\nu}):=\inf_{\hat{\gamma}\in\Pi(\hat{\mu},\hat{\nu})}\int_{\hat{\mathbb{G}}\times\hat{\mathbb{G}}}\hat{c}(x,y)\hat{\gamma}(\mathrm{d}x,\mathrm{d}y),$ (6)
then we have
$\mathrm{KT}(\mu,\nu):=\left(\mu({\mathbb{G}})+\nu({\mathbb{G}})\right)\left(\mathcal{W}_{\hat{c}}(\hat{\mu},\hat{\nu})-b\lambda\right)=\mathrm{ET}_{\lambda}(\mu,\nu).$ (7)
The proof is placed in Appendix §A.2.1.
Thus, we have reformulated EPT (4) for unbalanced measures as a corresponding standard complete OT (7) with nonnegative ground cost. Consequently, we bypass the technical challenges inherent in unbalanced settings and can leverage abundant existing results and approaches in the standard balanced setting for OT problems with unbalanced measures on a graph.
@@
Unlike existing approaches, e.g., as in [12, 36, 40, 72], the new ground cost $\hat{c}$ of the corresponding standard OT problem (7) for EPT is guaranteed to be nonnegative. Our calibration is essential for developing the associated OW from its standard OT as in [67, 32, 28, 17].
3.2 Orlicz-EPT
Following the approaches in [67, 32, 28, 17], we define Orlicz-EPT, which is EPT endowed with an Orlicz geometric structure, based on the standard OT problem (7) as follows: